OpenStudy (anonymous):

A cylindrical can, open at the top, is to hold 500 cm3 of liquid. Find the height and radius that minimizes the amount of material needed to manufacture the can.

OpenStudy (anonymous):

so far I got pi * r^2 * h = 500, h = 500/(pi * r^2).... Don't know the next steps from here.

OpenStudy (anonymous):

divide both sides by pi.

OpenStudy (anonymous):

leave r and h on one sides. Those are not constant and will vary

OpenStudy (anonymous):

And then I think you will need to use a derivative I just don't know eaxctly how

OpenStudy (anonymous):

K. So I'm still stuck. I got to thinking that V= pi * r^2 * h = 500 so h = 500/(pi*r^2) Surface Area = pi*r^2 + 2*pi*r*h. then yeah I think I do need to take a derivative or something

OpenStudy (anonymous):

its optimization. you need to take the volume equation and take the derivative and plug it into your objective equation

OpenStudy (anonymous):

@ChukRock Which is the surface area of the object?

OpenStudy (anonymous):

surface area = 2(pi)r(h)+2(pi)r

OpenStudy (anonymous):

surface area = 2(pi)r(h)+(pi)r^2 I think it's this one.

OpenStudy (anonymous):

u sure I need to take the derivative of the volume equation ? with the two variables and h and not the derivative of the surface area with h = 500/(pi*r^2) ?

OpenStudy (anonymous):

to be honest i'm not positive cause i dont have my notes with me